OpenStudy (anonymous):

Find the intercepts of the graph of the equation. x^2-9/x^2-4

OpenStudy (amistre64):

when x=0, what does y equal? when y=0, what does x need to equal? hint, the top goes zero

OpenStudy (anonymous):

So I do 9/4?

OpenStudy (amistre64):

that will be the y intercept, yes

OpenStudy (amistre64):

and recall that in order for a fraction to be zero; then the top part has to be zero ... as long as the bottom part doesnt zero out as well

OpenStudy (anonymous):

2.25

OpenStudy (amistre64):

for the y intercept, yes. 2.25

OpenStudy (anonymous):

I think I am doing it wrong...

OpenStudy (anonymous):

so (0,2.25)

OpenStudy (amistre64):

y = x^2-9/x^2-4 9/4 = 0^2-9/0^2-4 = 2.25 y = x^2-9/x^2-4 0 = 0/x^2-4, so for this we will need to determine x^2-9 = 0

OpenStudy (anonymous):

What do I plug in?

OpenStudy (amistre64):

umm, for the second one? you plug in y=0 and solve for x

OpenStudy (amistre64):

x^2 - 9 = 0 , when x=?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

lets see: 0^2 - 9 not= 0 , so x=0 is not correct

OpenStudy (anonymous):

-9

OpenStudy (amistre64):

ummm, lets see (-9)^2 - 9 = 0 81 - 9 ... not equal 0, so x=-9 is not it

OpenStudy (anonymous):

A plot may help.

OpenStudy (amistre64):

x^2 - 9 = 0 , lets start by adding 9 to each side x^2 -9 +9 = 0 + 9 x^2 = 9 , lets take the square root of each side to undo that ^2 sqrt(x^2) = sqrt(9) , lets take the square root of each side to undo that ^2 x = +- sqrt(9)

OpenStudy (amistre64):

and since: sqrt(9)^2 - 4 not equal 0, the denominator is safe .....

OpenStudy (anonymous):

x= to 0 though

OpenStudy (amistre64):

not sure i follow you on that ine

OpenStudy (anonymous):

I am confused sorry

OpenStudy (anonymous):

what do you mean the denominator is safe?

OpenStudy (amistre64):

we are not allowed to divide by zero in math .... so as long as a solution for the top part does not affect the bottom part of the fraction, we are good to go.

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

\[\frac{x^2-9}{x^2-4}\frac{\text{<-- needs to be zero}}{\text{<-- cant be 0}}\]

OpenStudy (amistre64):

\[3^2-9=0\]\[3^2-4\ne0\]safe \[(-3)^2-9=0\]\[(-3)^2-4\ne0\]safe

OpenStudy (anonymous):

So how do I find the intercepts?

OpenStudy (anonymous):

so y=3?

OpenStudy (anonymous):

or x=3

OpenStudy (amistre64):

.... to find the y intercept, make x=0: (0,y) to find the x intercept (may be many of them), make y=0: \((x_1,0),(x_2,0),...\)

OpenStudy (anonymous):

(0,3)

OpenStudy (amistre64):

\[y=\frac{x^2-9}{x^2-4}\] yint is x=0 \[y=\frac{0^2-9}{0^2-4}=\frac94~;~yint~(0,\frac94)\] xint is when y=0 \[0=\frac{x^2-9}{x^2-4}\]since a fraction is 0 when the top is zero ...\[0=\frac{0}{x^2-4}~;~(x^2-9=0)~;~(x^2-4)\ne0\]

OpenStudy (amistre64):

xints are therefore (3,0) and (-3,0)

OpenStudy (anonymous):

So they basically have to equal

OpenStudy (amistre64):

no, thats just a coincidence of this particular setup. It not a general conclusion

OpenStudy (amistre64):

if im reading you correctly that is ...

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

good luck